OpenStudy (anonymous):

Wood for a fireplace is often sold by the cord. A cord is 8 ft by 4 ft by 4 ft. How many cubic feet are in a cord of wood?

OpenStudy (mendicant_bias):

Okay. So a cubic foot is simply 1 foot by 1 foot by 1 foot, or ft^3. You can find out the number of cubic feet by multiplying all of the given dimensions by one another to give you volume, because the typical formula for cubes or rectangular prisms for Volume is such: \[v = l*w*h\] So, for this:\[V = 8*4*4 = ?\] Does this help?

OpenStudy (mendicant_bias):

And this makes sense in terms of units, because you have \[(8 ft.)(4 ft.)(4 ft.) = (32ft.^{2})(4ft.) = 128 ft.^{3}\]

OpenStudy (anonymous):

yes thanks

OpenStudy (mendicant_bias):

You're welcome.